Instrument Log 

SAS-HE          C1                  CF Lamont, OK

Date

Technician

Comments

7/22/2016

MS

A pop up window that said SASpect occurred causing the program to stop.  At 1301 GMT closed out pop up window and program automatically started back up.  RTS at 1316 GMT.

7/20/2016

MS

Instrument had stopped running at 1713 GMT due to a Relative Humidity pop up window.  Closed out pop up window at 1713 GMT and the instrument automatically started back up.  RTS 1728 GMT.

7/19/2016

MS

Noticed that the SASpect program had stopped.  A pop up window came up saying that the SASpect has stopped working.  A problem caused the program to stop working correctly.  Windows will the program and notify you if a solution is available.  Closed out the SASpect window and restarted the instruments program at 1934 GMT.  RTS at 1949 GMT.

7/18/2016

MS

Instrument went to sleep at 00:00 GMT on 7-16-2016 and never woke back up.  Instrument goes to sleep at 0500-1100 GMT every day.  Woke instrument back up by closing out pop up window at 1259 GMT, 7-18-2016.  RTS at 1314.

07/13/2016

MS

The diffuser was not getting fully shaded, so at 2025 GMT adjusted and tightened the motor belt that drives the shadow band, then re-aligned the instrument. RTS at 2050 GMT.

7/7/2016

MS

EDIT SAZ pop up window has occurred again causing program to stop running.  Closed out the pop up window at 1322 GMT,  program started back up automatically.  RTS 1337 GMT.

7/5/2016

MS

A EDIT SAZ pop up window has occurred causing instrument to stop as of 7/2/16 at 0100 GMT.    Closed out pop up window at 1308 GMT,  and program automatically started back up.  RTS at 1323 GMT.

7/1/16

KT

SDS reported no data was collected from the instrument. At 1224 GMT during computer checks, there was an Edit Collector window that was open on the desktop that caused the program to quit running. Closed the Edit Collector window and the program began normal operations. System is online and operational at this time. RTS at 1239 GMT.

6/29/16

Mentor (Albert)

Looked over the SAS-He and found the shadow band to be out of alignment with respect to the hard stops. It appeared the pulley slipped on the shaft of the hard stop. Realigned the pulley, calibrated the inner band & azimuth, then restarted software and all appears to be working.  RTS at 1600 GMT.

06/20/2016

MS

Diffuser is not getting fully shaded by shadow band.  Followed instructions from manual and still could not prevail.  Contacted mentor and will proceed 06-21-2016 to get instrument to fully shade again with new procedures.  RTS

6/16/2016

MS

Noticed during PM rounds this morning that the diffuser was not shading correctly on the CW and the CCW band rotation.  Used the manual to properly realign the shadow band so that it shades both the CW and the CCW bands.  Got it to shade only on one of the bands.  Mentor has been contacted about issue and he is trouble shooting the issue.  RTS 1815 GMT.

5/25/16

KT

SDS reported no data collected from the instrument. At 1219 GMT found Edit SolEl window was open on the desktop which caused the program to lock up. At 1223 GMT closed the Edit SolEl window and the program began normal operations. RTS at 1234 GMT.

5/23/16

KT

DS View page showed that the SAS-He instrument was red. At 1214 GMT found Auto Scale window opened on the desktop which caused the program to quit running. Closed the Auto Scale window and the program began normal operations. RTS at 1229 GMT.

5/16/16

KT

The program was not running. At 1337 GMT rebooted the computer. At 1341 GMT the computer came online and began normal operations. RTS at 1349 GMT.

5/11/2016

MS

Program had stopped running due to a Edit Relative Humidity issue.  Closed out the pop up window and program started back up on its own.  RTS at 1319 GMT.

4/22/2016

MS

Program had stopped running due to a Edit Relative Humidity issue.  Restarted program at 1313 GMT, RTS at 1527 GMT.

4/11/2016

MS

Instrument had stopped collecting and no data had been collected.  A window was up that said SASpect has stopped working.  Closed out program and restarted at 1947 GMT,  RTS 2002 GMT.

4/11/2016

MS

No data had been collected from SAS-He since 4-9-2016.  Remote status check found a Humidity edit box open and SASpect program was not updating.  Closed edit box and RTS at 1322 GMT.

3/28/16

CAM

No data had been collected from SAS-He since 3/26.  Remote status check found a Humidity edit box open and SASpect program was not updating.  Closed edit box and RTS at 0208 GMT.

3/4/2016

MS

Noticed that the program had stopped running due to a pop up window that said, EDIT Relative Humidity.  Instrument had stopped at 1300 GMT.  Closed out pop up window and instrument automatically started back up at 1429.  RTS at 1444 GMT.

3/3/2016

MS

Checking the VM's this morning at 1412 GMT,  I noticed that a pop up window that said, EDIT TILT had stopped the program from running at 1300 GMT.  Closed out the EDIT TILT pop up window and the instrument program automatically started back up.  RTS 1427 GMT.

2/19/2016

MS

Noticed during PM rounds that the instrument program had stopped running at 1501 GMT, 2/19/2016.  Started instrument program back up at 1501 GMT.  RTS at 1516 GMT, 2/19/2016.

2/11/2016

MS

A pop up wind was displayed, it said EDIT TILT.  This caused instrument to stop running.  Closed out EDIT TILT pop up  window and the program automatically started back up.  RTS 1548,  2/11/2016.

2/11/16

JEM

The new air compressor for the dry air system has been received and needs to be installed.  Installed the new air compressor from 2100 GMT to 0030 GMT on 2/12/16 GMT.  Leak checked the system and verified OK.  System was working as expected and RH values were low.

2/2/16

KT

The Air compressor belt had broken which provides air flow to the instrument keeping the RH value down. At 1630 GMT there were no replacement belts available to repair the compressor. James Martin and Matt Gibson located a spare air compressor that was installed while a new main Air Compressor could be ordered. As of now the instrument has air flow and is keeping the RH value down to normal values. RTS at 1645 GMT.

2/1/16

KT

Instrument mentor stated that the RH value for the instrument was steadily on the rise. At 2105 GMT found the belt for the compressor that keeps air flowing through the instrument had broken. Troubleshooting still in progress to locate a replacement belt for the air compressor. RTS at 2130 GMT.

2/1/16

KT

Instrument Mentor reported puzzling measurements coming from the instrument and requested help to perform testing. At 2000 GMT placed a 2"x4" aluminum foil square to cover the diffuser. At 2010 GMT the aluminum foil was removed. RTS at 2015 GMT.

1/12/2015

MS

Noticed that the instrument program had stopped running around 1100 GMT 1/12/2016.  Started instrument program back up at 1422 GMT,  RTS at 1437 GMT 1/12/2016.

12/24/15

MS

Noticed that the SAS-He Full Control was closed out and program was not running.  Restarted SAS-He Full Control program at 1405 GMT. RTS at 1420 GMT.

12/15/15

MS

Computer needed to be restarted following scheduled software and security patch updates.  At 1404 GMT computers were updated and rebooted by the IT department, some data might be lost.

12/14/15

MS

Programed had stopped on 12/13/2015 at 2100 GMT.  Started program at 1421 GMT.  RTS at 1436 GMT.

12/11/15

MS

Program was not running during PM rounds, may have lost 12 hours of data.  Stopped running at 0200 GMT, 12/11/2015.  Restarted program but got the same pop up windows, took screen shots of pop up windows and contacted mentor.  As of 1500 GMT 12/11/2015 mentor has the problem fixed.  RTS 1500 GMT.

12/8/15

MS

Program was not running and noticed that it was closed out as well.  A pop up window was up on the screen and was labeled NI USB-6229 "Dev7"-NI Device Monitor.  Three options were available,  Configure and test this device, Test this device, Do nothing.  Closed out pop up window without choosing any of the three options. Oened the program and started instrument back up.  RTS 1419 GMT.

12/7/15

KT

SDS stated no data collected from the instrument. At 1447 GMT upon inspection, there was a USB Device detected window that was up on the home screen which caused the program to quit running. At 1449 GMT closed the notification window and the program began normal operations. RTS at 1502 GMT.

11/16/15

MS

Program for the SAS-He had stopped running around 0200 GMT 11/14/2015.  Closed out program on instrument and restarted program at 1433 GMT 11/16/2015.  RTS 1433 GMT.

11/23/15

KT

SDS reported not data collected from the instrument. During inspection error window caused the program to lock up. After clearing out the error window the program was restarted and just asked for control, Alt and delete to restart the computer. Attempted a couple of more times. Called and left message for instrument mentor. Tim Grove pulled the computer to rebuild the hard drive. RTS at 1404 GMT.

11/20/15

KT

SDS stated no data collected from the instrument. Found message on screen stating - National Instrument.DAQmx.DAQ Exception.error code could not be found. At 1444 GMT closed the window and the program began normal operations. RTS at 1459 GMT.

11/18/15

KT

SDS stated no data collected from the instrument. At 1350 GMT observed that there was an Edit Collector window that was open that has caused the program to quit running. At 1355 GMT closed the window and the system began normal operations. RTS at 1405 GMT.

11/2/2015

KT

SDS reported no files from the instrument. At 1435 GMT Edit NIR Ad Temp window was displayed causing the program to quit running. Closed the window out and the program began normal operations. RTS at 1450 GMT.

10/19/2015

MS

The SAS-He computer program had stopped running, due to locking up on 10/17/2015  0300 GMT.  At 1319 GMT the program window was closed and the computer was rebooted.  At 1334 GMT the computer came back online and began normal operations.  RTS 1334 GMT.

10/5/15

KT

SDS reported no files from the instrument. At 1334 GMT found the program locked up and not running. At 1338 GMT closed the program window and restarted the computer. At 1340 GMT the computer came online and the program began normal operations. RTS at 1349 GMT.

9/28/15

KT

SDS reported no data from the instrument. At 1224 GMT observed that the program was locked up. At 1228 GMT cleared the error window and started the program. System is online and operational at this time. RTS at 1239 GMT.

9/20/15

CAM

No data had been collected since 2308 GMT on 9/18.   Remotes status check found EDIT TEMP dialog box open.  Closed dialog box and program resumed at 1751 GMT.

9/18/15

KT

The program had locked up. At 1305 GMT closed the window and started the program. At 1315 GMT system was online and operational. RTS at 1320 GMT.

9/8/15

KT

The computer program was not running. At 1225 GMT started the program. At 1230 GMT instrument came online and began normal operations. RTS at 1240 GMT.

9/4/15

KT

The computer program was not running. At 1335 GMT restarted the program. At 1339 GMT the computer came online and began normal operations. RTS at 1350 GMT.

8/27/15

CAM

No data had been collected from SAS-He since 1908 GMT on 8/26.  Logged in remotely and found NationalInstruments.DAQmx.Daq.DaqException error displayed. Error stated: Attempted to read samples that are no longer available.  The requested sample was previously available, but has since been overwritten.  Time displayed in SASpect status window was 1754 GMT, 8/26.  OK’d error message and SASpect program began updating.  Closed SASpect and restarted.  RTS at 1336 GMT.

8/26/15

KT

The SAS-He computer program had stopped running, due to locking up. At 1455 GMT the program window was closed and the computer was rebooted. At 1500 GMT the computer came online and began normal operations. RTS at 1510 GMT.

8/19/2015

MS

Program on the SAS-He had stopped running at 1900 GMT 8/18/2015.  Restarted program on the SAS-He at 1303 GMT 8/19/2015, RTS at 1318 GMT 8/19/2015.

8/6/15

KT

Computer program error caused the SAS-He program to stop. At 1518 GMT closed the error and restarted the program. At 1523 GMT selected Run Continuous and the instrument and computer began normal. RTS at 1533 GMT.

7/30/15

KT

The computer program was not running. At 1440 GMT restarted the program. RTS at 1455 GMT.

7/27/2015

MS

Program and graphs had stopped running and updating at 1400 GMT 7/25/2015.  Restarted computer at 1300 GMT 7/27/2015. RTS 1315 GMT 7/27/2015.

7/21/2015

MS

Program had stopped collecting data 0100 GMT 7/21/2015.  Restarted program at 1304 GMT, RTS at 1319 GMT.

7/20/2015

MS

Program was frozen when found at 1255 GMT on daily PM rounds.  Restarted SASpec Version program and clicked on Run Continuously.  RTS 1310 GMT.

6/30/2015

MS

Instrument program had stopped running at 0700 GMT 6/30/2015.  Started program back up at 1311 GMT.  RTS at 1326 GMT.

6/29/2015

MS

Failed FTP.  The Intensity  Wavelength graphs were not graphing and the shadow band arm was not shading the diffuser fully.  At 1450 GMT, realigned shadow band and restarted the program at 1530 GMT.  RTS 1530 GMT.

6/26/16

CAM

No data had been collected from SAS-He since 0308 GMT. Checked status remotely and found EDIT GPS SYCN window open.  SASpect program was also open, but not updating and no graphs were displayed.  Closed GPS SYNC window and restarted SASpect.  RTS at 1327 GMT.

6/18/15

JEM

The SAS-He freezer thermostat has been losing its settings every time the power goes out.  Checked the thermostat batteries and they were a little low and there was some battery acid corrosion on one of the terminals.  Removed the corrosion, cleaned the terminal and installed a fresh pair of batteries.  RTS @1445 GMT.

6/18/15

MS

SAS-He program stopped working at 0400 GMT.  Restarted SAS-He program at 1323 GMT.  RTS at 1338 GMT.

6/17/15

JEM

There was a power outage at the Central Facility that reset the SAS-He freezer thermostat.  Internal freezer temperature was up to 15C.  Reset the thermostat to 35F (1.6C).  The freezer began to operate.  Internal temperature should return to normal within an hour.  RTS @1400 GMT.

6/10/15

MS

Thermostat was off for the SAS-He chiller at 1853 GMT.  Turned on thermostat for the SAS-He chiller at 1853. RTS at 1908 GMT.

6/10/15

CAM

Power to all ground mounted transformers is being cut in order to connect the new power panel near the AOS trailer and run power to new PGS building.  Instrument needs to be shut down.  Shut SAS-He down at 1352 GMT.  Restarted once power was restored and RTS at 1544 GMT.

6/9/15

MS

SAS-He program had stopped and was not collecting data at 0700 GMT on 6/9/2015.  Restarted SAS-He Version program at 1327 GMT.   RTS at 1342 GMT.

6/2/15

JEM

The air compressor for the dry air system has started tripping the breaker again.  The mentor is currently on site working on upgrades to the instrument so the compressor problem is not affecting any data.  Troubleshooting determined that the compressor, regulator and start switch were not the cause of the problem and that the new motor is failing.  R/R the compressor motor and RTS the dry air system at 2000 GMT.

6/2/15

MS

SAS-He is shut down.  Mentor is on site and troubleshooting the SAS-He.

5/29/15

CAM

Chiller RH is running high for both SAS-He and Sas-ZE.  T/S found the breaker was tripped for the dry air system compressor.  The breaker was also found tripped yesterday.  Reset breaker and compressor kicked on normally.  Let tank pressurize to 150psi and then manually drained tank to 100psi to force compressor back on. Compressor again operated normally.  Repeated this once more with no problem.  RTS at 1610 GMT.  Will continue to monitor status.

4/21/15

MS

SAS-He was not following the sun at 1447 GMT.  Program had stopped running.  Restarted computer at 1502 GMT and started program back up.  RTS 1502 GMT.

3/18/15

PD

SASpect program is up but no graphs are updating.  Exited program and restarted.  Graphs are now shown and updating.  RTS at 1459 GMT.

1/23/15

JEM

The air compressor was found inoperable during morning rounds.  R/R the air compressor motor with a newly purchased motor.  RH values began to drop after the compressor was restarted.  RTS @1900 GMT.

1/23/15

KT

The Purge Air Pressure was reading 0. The main outside Air Compressor was off. At 1512 GMT checked the breaker. Breaker didn't appear to be tripped but reset the breaker. No change in status, the Air Compressor did not come on. Pressed the reset switch on the compressor and reset the breaker again. No response from the Air Compressor. Trouble shooting in progress. RTS at 1530 GMT.

1/20/15

KT

The computer program was not running. At 1550 GMT restarted the program. At 1557 GMT system was online and operational. RTS at 1605 GMT.

1/22/15

CAM

Chiller RH reading over 40%.  Checked air flow from dry air system and found it was weak.  Increased air flow at 2030 GMT and RH started dropping initially, but began to rise again after several minutes.  Checked for leaks around chiller seal but found none.   Confirmed good air flow from line inside chiller.  Increased the flow slightly more and planned to watch overnight.  The following day, the dry air compressor motor had failed.  Motor has been experiencing problems for some time and a replacement was recently ordered and received.  Technicians in process of replacing motor and will continue troubleshooting following compressor maintenance.

1/13/15

KT

The instrument was locked up. At 1501 GMT stopped the program and rebooted the computer. At 1506 GMT the computer came online and the instrument rotated to it's correct position. System is online and operational. RTS at 1516 GMT.

1/12/15

JEM

Ken Teske notified me that on morning rounds he found the air compressor inoperable again.  Troubleshooting found the start capacitor had failed again.  Due to continued issues with the air compressor motor, I ordered 2 new motors, one for use and one for a spare.  Was able to find another capacitor that was close enough to the correct size to get the motor running again.  Will install the new motor either when it arrives or when the current motor fails again, whichever happens last.  RTS @2000 GMT.

1/12/15

KT

The main air compressor that provides air for the Dry Air system was not operating. During inspection the breaker for the main air compressor was tripped. Attempted to reset the breaker twice and still could not get the air compressor to return to service. Trouble shooting is in progress. RTS at 1717 GMT.

1/12/15

KT

The instrument needed to be aligned. At 1643 GMT adjusted the inner band as well as the Azimuth. At 1656 GMT after alignment was completed the instrument was placed back into operation. RTS at 1700 GMT.

12/9/14

JEM

The Chiller RH began climbing again approximately 1 hour after the air compressor was returned to service on 12/8.  Tried resetting the breaker and again it would trip after a few seconds of operation.  Disassembled the motor and checked the internal centrifugal start switch again and it was not stuck and was working properly.  Suspect a problem with the motor.  R/R the motor with the spare.  Reset the breaker after installing and connecting the spare motor.  The breaker would immediately trip and the motor would not even try to run.  Removed the motor and took both motors to the repair lab for further inspection.  The spare motor would also trip the breaker in the repair lab and further inspection found a burnt spot in one of the windings causing a short.  Powered on the original motor in the lab and it ran just fine and current draw was only 7 amps.  Returned the original motor to the optical trailer air compressor  and again it would trip the breaker after a few seconds of operation.  Current draw was measured at ~40 amps.  Inspection of the cut in/out switch connected to the high pressure cutoff revealed that all of the contact pads had a large amount of carbon buildup and damage from arcing.  R/R the cut in/out switch with a new one.  Due to all of the times the breaker has tripped in the past the breaker was also replaced with a new one.  Powered the compressor on at 2200 and it ran normally.  Dry air flow returned to the SAS-He and the RH began to drop.  RTS @2200 GMT.

12/8/14

JEM

Received an email from the mentor that the RH in the SASHe Chiller began to climb on 12/5 and was currently very high.  Checked the air compress for the dry air system and it was not operating.  Tried resetting the breaker several times and the breaker would trip after a few seconds of operation.  Disassembled the air compressor motor and the internal centrifugal switch was stuck in the on position, not allowing the start winding to disengage once the motor was up to speed causing excessive current draw.  Replaced the internal centrifugal start switch with a new one.  Reassembled the motor and powered on the compressor.  Dry air flow returned to the SASHe and the RH began to drop.  RTS @1830 GMT.

12/1/14

JM/CM

The dry air compressor has tripped the breaker several times in the past few weeks.  Replaced the start capacitor for the compressor motor at 1900 GMT.  One of the terminals broke off the original capacitor when it was disconnected, so it may have been the source of the problem.

12/1/14

KT

The Outside Air Compressor was not running. At 1520 GMT found that breaker had been tripped. Reset the breaker and the Air Compressor powered back on. RTS at 1535 GMT.

11/12/14

JEM

Was notified by site ops that the air compressor used to supply dry air to the freezer is not working. The circuit breaker was found tripped and when trying to reset the breaker it trips again after about 1 second.  Troubleshooting found that the start capacitor on the motor has failed.  R/R the start capacitor and reset the circuit breaker.  Compressor began working normally.  After the compressor pressured up, observed the freezer humidity begin to drop.  RTS @1830 GMT.

11/10/14

MS

Shadow band was not stopping to shade the sensor, noticed at 1543 GMT.              At 1601 GMT rebooted computer and then re-adjusted shadow band through the computer, shadow band is now shading sensor.

11/7/14

MS

GPS and Motor boxes were red at 1521 GMT, shadow band was not aligned or shading properly.  Stopped and restarted program at 1558 GMT, realigned shadow band with computer and RTS at 1618 GMT.

10/1/14

CAM

The dry air system compressor motor failed on 9/30.  Removed the failed motor and replaced with a spare rebuilt motor.  Removed smaller compressor that had been temporarily plumbed in yesterday and placed the original compressor back in service at 2000 GMT.

9/30/14

CAM

Technician found during daily rounds that the compressor motor for the dry air system had failed.  Temporarily connected a smaller compressor to the dry air system until the large compressor can be repaired.  RTS at 2115 GMT.

8/14/14

CAM

The RH inside the SAS-He refrigerator has risen over the past several weeks and is at times near 60%.  Found that SAS-He, SAS-Ze, SORTI, and SWS were all on the same regulated dry air line.  There was another regulated air line that had only RSS on it.  Reconfigured the air lines so that SAS-He and SORT were on a line by themselves.  RH dropped to near 0% for both SAS instruments and appears to be holding.  RTS at 2130 GMT.

5/29/14

Mentor

Flynn   

Instrument mentor on site.  Found a few teeth were missing on the radial drive gear, which was causing the azimuth tracking to be off.  Mentor pulled instrument to be sent in for repair.

3/7/14

KT

The SAS-He alignment needed to be adjusted. At 1430 GMT adjusted the shadow band as well as the Azimuth. At 1425 GMT started the program and the instrument began normal operations. RTS at 1454 GMT.

3/4/14

KT

System shut down due to power outage that occurred on 3/2/2014 at 2120 GMT. CF was closed 3/3/2014 due to snow storm. CF opened on 3/4/2014. Powered on the computer at 1350 GMT. System came online at 1353 GMT. RTS at 1405 GMT.

1/30/14

KT

Kay Electric had arrived at the CF to shut power down to perform maintenance. Time of power outage was from 1614 GMT to 1623 GMT. Once power was restored the SAS-He instrument needed to be realigned. At 1730 GMT made adjustments to align instrument. RTS at 1830 GMT.

1/14/14

CAM

Instrument mentor advised that the refrigerator wasn’t cooling and suspected the controller may have been accidentally shut off during previous maintenance visit on 12/20/13.  Powered temperature controller on at 2000 GMT.  RTS status O.K.

9/23/13

KT

The shading needed to be adjusted on the instrument. At 1630 stopped the program and adjusted the inner band. RTS at 1645 GMT.

7/3/13

KT

The wavelength graphs on the display were not normal for the instrument. Called and talked to instrument mentor and began trouble shooting. The Fiber optic line that connects to the diffuser on the instrument was not connected. Connected the fiber optic line to the diffuser and started the program. Instrument is online and operational. RTS at 1800 GMT.

6/17/13

MS

The SAS-He program was not running.  At 1412 GMT rebooted the computer. RTS at 1427 GMT.

5/31/13

KT

The inner band and Azimuth needed to be adjusted. At 1420 GMT aligned the inner band and also adjusted the Azimuth. Instrument is online and operational. RTS at 1444 GMT.

5/8/13

CAM/JEM

Instrument was removed on 4/17 so the wooden decking and stairs could be replaced with metal.  Construction work is complete and instrument needs to be reinstalled.  Reinstalled SAS-HE, powered on and started program at 1524 GMT.  Aligned shadow band and RTS at 1600 GMT.

4/17/13

CAM/JEM

The optical trailer wooden decking and stairs are to be removed and replaced with metal decking and stairs.  Need to remove all instruments so removal of old decking can start tomorrow.  Stopped SASpect program at 1401 GMT.  Disconnected fiber, data and power cables at the computer end and pulled cables out through the wall.  Removed SAS-He instrument and stored inside trailer with coiled cables.  Removed mounting hardware (Unistrut) and stored in S/R warehouse.  Instrument will remain offline until new deck and stairs are complete. 

4/12/13

KT

The inner band and Azimuth needed to be adjusted. At 1330 GMT made adjustments to the inner band and Azimuth. RTS at 1400 GMT.

3/6/13

PAD

Azimuth is not in correct alignment.  Stopped SASpect program and zeroed inner band.  Calibrated inner band offset.  Calibrated azimuth offset.  Experiencing trouble calibrating both inner band and azimuth offsets.  Inner band will not allow movement of band to any negative number entered.  Number changes to 0.000 after send command is selected.  Azimuth calibration will change any number entered >100 to 100.000 after send command is selected.  Both finally accepted entered numbers after numerous attempts.  Instrument is now shading correctly.  Notified instrument Mentor.  RTS at 1634 GMT.

2/27/13

PAD

The site lost power at approximately 0350 GMT on 2/26 due to a severe winter storm.  Power was restored at 1920 GMT.  On 2/27/13, observed shading was not correct.  Adjusted azimuth and inner band offsets at 1828 GMT.  RTS at 1843 GMT.

1/23/13

PAD

Shadowband is not fully shading diffuser.  Stopped SASpect program and zeroed inner band.  Changed inner band offset as needed to rezero.  Restarted program.  RTS at 1658 GMT.

1/14/13

CAM

After reinstalling the SWS, noticed the dry air system was not providing air.  T/S found the copper air line between the compressor and the building was plugged with ice.  Disconnected copper line at building entrance, thawed the ice, and cleared water from line.  Reconnected line and confirmed air flow.  RTS at 2030 GMT.   SAS-He, RSS and SORTI were also affected by lack of air.

12/13/12

KT

The computer program was not running. At 1450 GMT started the program. RTS at 1505 GMT.

11/27/12

KT

The instrument needed to be realigned. Chris Martin and I realigned the SAS-He. Left voicemail with instrument mentor notifying him of the alignment issue. RTS at 1440 GMT.

11/20/12

CAM

Local electrical service was shut off at 1410 GMT so that Kay Electric could upgrade its power lines from 7.2 KV to 14.4 KV in Grant County.  SASHE was gracefully shut down at 1341 GMT in anticipation of power outage.  Power was restored at 2115 GMT after electrical upgrades were complete.  Restarted SASHE and RTS at 2146 GMT.

11/13/12

CAM

Site power was partially shut down at 1400 GMT in order to replace the 7 transformers that service the buildings and all instruments south of the main compound.  Power was restored at approximately 1930 GMT after all transformers were replaced.  Restarted SASHE at 2107 GMT.  Aligned band and RTS at 2151 GMT.

9/20/12

PAD

Azimuth position is not in correct alignment with sun.  Site experienced power outage on 9/19/12 and SASpect program was started remotely at 1952 GMT.  Reset azimuth offset.  Offset value changed slightly from 16.11 to 16.31.  Misalignment most likely due to instrument losing where zero was following power outage on 9/19/12.  RTS at 1444 GMT.

9/19/12

PAD

CF experienced a power outage between 1528-1535 GMT.  Remotely accessed SAS-He computer and error message displayed:  National Instruments DAQmx. Daq Exception.  Device cannot be accessed.  Cleared error message and restarted SASpect program.  RTS at 2007 GMT.

6/5/12

PAD

Inner band is not shading diffuser.  Band is coming further down on one horizon than the other.  Suspect mechanical limit stop had slipped on shaft.  Opened instrument enclosure and found mechanical stop set screw was tight.  Adjusted inner band to come down evenly to horizon in both directions.  Adjusted inner band offset and restarted program.  Inner band is now shading correctly.  Suspect bird may have landed on inner band causing it to slip on shaft and get out of alignment.  RTS at 1507 GMT. 

5/16/12

PAD

Inner band offset has needed adjustment several times over the last couple weeks.  Inner band shaft was recently lubricated but that did not correct improper shading issue.  At 1929 GMT, stopped measurement and opened instrument side panel.  Found inner band mechanical stop set screw was loose.  Tightened set screw and reset inner band offset.  Started program.  Shadowband is now shading correctly.  RTS at 1954 GMT.

5/14/12

PAD

Shadowband is not shading sensor.  Adjusted inner band offset and saved configuration.  Started SASpect program and shadowband is now shading correctly.  RTS at 1455 GMT.

5/10/12

PAD/KT

SAS-He shadowband is not shading correctly.  Zeroed inner band and it is not in correct position for zero degrees.  Initialized calibration offset for inner band.  Inner band would not move to any negative position entered.  When negative position was entered and command was sent, entered number would change to zero.  Checked lower limit in properties and it was -100 degrees.  Inner band finally moved to negative position after third attempt.  Set offset for inner band, saved configuration and started measurements.  Shadowband is now shading correctly.  RTS at 1622 GMT.

5/8/12

PAD

Notified by mentor that data indicates shadowband is not always fully shading diffuser.  Previous troubleshooting has found that Inner Band shaft needs lubricated occasionally or it will drag enough to cause shadowband alignment to be incorrect.   Stopped measurements at 1435 GMT.  Removed top cap on drive end of shadowband.  Lubricated shaft with greaseless lubricant.  Installed cap and reset azimuth and inner band offsets.  Verified shadowband was shading correctly.  RTS at 1525 GMT.

4/30/12

PAD

Error is shown:  Saspect has encountered a problem and needs to close.  GPS sync status box is red.  Motor error is flashing red.  Cleared error and rebooted computer.  At 1512 GMT, found shadowband was not shading correctly.  Changed inner band offset from 0 to -2.988.  Started measurements but shadowband is still not completely shading diffuser.  Adjusted azimuth offset from 17.85 to 16.15.  Started measurements and shadowband is now shading correctly.  RTS at 1542 GMT.

4/26/12

PAD

Shadowband is not completely shading diffuser.  Stopped program and checked inner band zero.  Adjusted azimuth offset from 15.8 to 17.8.  Started program and shadowband is now completely shading diffuser.  RTS at 1635 GMT.

4/23/12

PAD

Shadowband is not shading diffuser.  Stopped SASpect program and zeroed inner band.  Inner band appeared to be at zero.  Restarted program but shadowband is still not shading correctly.  Stopped program and reset inner band offset.  Offset was changed from -2.9... to 0.0.  Saved configuration and restarted program.  Shadowband is now shading correctly.  RTS at 1534 GMT.

4/9/12

PAD

Notified by mentor data taken over the weekend shows shadowband is not completely shading diffuser and needs aligned.  It was overcast this morning and shading could not be checked.  At 1722 GMT, stopped measurement and adjusted shadowband offset.  New offset is -2.98...  Saved configuration and restarted measurements.  Shadowband is now shading correctly.  RTS at 1737 GMT.

3/12/12

PAD

Last SDS collection was on 3/11/12 at 0900 GMT.  Remotely accessed computer and error was shown:  Intel PROSet Wireless Out of Memory.  Cleared error and SASpect error was shown:  SASpect.exe Application Error.  Application has generated an exception that could not be handled.  Process ID=Ox108(264.  Thread ID=Ox1620(5664).  Ok to terminate.  Cancel to debug.  Selected Ok to end program.  Rebooted computer and placed program in continuous experiment.  Notified computer department who then disabled wireless adapter.  RTS at 1230 GMT.

2/29/12

C. Flynn

Mentor replaced both motors and ribbon cable late in the afternoon.  Azimuth and shadow band are now working as they should.

2/23/12

PAD

SASpect program indicates shadowband is moving but it is not.  Stopped program.  Shadowband will not respond to commands.  Saved configuration.  Shadowband will now respond.  Started program but shadowband is not shading correctly.  Set offset for azimuth and inner band.  Started program.  Shadowband is now shading correctly.  RTS at 1732 GMT.

2/22/12

PAD

SAS-He shadowband has not been shading correctly and repeatedly giving motor errors.  Mentor requests inner band motor replaced with spare provided.  Shut system down at 1904 GMT and R/R inner band motor.  After install, tried to move inner band.  Inner band would move very slowly but position would increase quickly up to 600 deg when inner band had only moved 20 deg.  Contacted mentor.  At 2231 GMT, worked with mentor via phone.  Inner band was acting as before.  Mentor saved configuration and then inner band would move correctly.  Zeroed shadowband and ran single experiment.  Shadowband was then not responsive but no errors were given.  Mentor moved azimuth to zero and then shadowband would move.  Mentor started continuous experiment and instrument moved into correct azimuth position and shadowband is now shading correctly.  Old motor model MD13CRD17A4-EQ, SN 036100206.  New model MD13CRD17A4-EQ, SN 036100214.  RTS at 2231 GMT

2/8/12

PAD

Last SDS collection was at 0100 GMT.  Saspect program is shown but covered by screen saver.  Used task manager to close program.  Rebooted computer and started continuous measurements.  Motor errors are still being given after restart.  RTS at 1611 GMT.

1/25/12

PAD

Last SDS collection was at 0200 GMT.  SASpect program is not running and error is shown:  Intel PROSet Wir...  Out of memory.  Cleared error and tried to start SASpect program.  Computer would not respond to icons or commands.  Tried to restart computer but it would not respond.  Power cycled computer, started program, zeroed inner band and azimuth.  Motor errors are shown after restart.  Motor errors are an ongoing problem mentor is aware of.  Instrument moved into correct azimuth position but shadowband is not shading or doing what instrument status shows.  RTS at 1635 GMT.

1/24/12

PAD

SAS-He recently has shown Az stall errors and Motor errors.  Mentor advises to remove part of the material used to seal azimuth cover plates to bottom of instrument.  Suspects change in temperature and other factors are causing too much friction between cover plates and instrument causing azimuth stall errors.  Stopped program and removed azimuth cover plates at 1955 GMT.  Removed inner half of material on both plates and reinstalled plates.  Opened instrument and checked connectors on drive motors.  No problems found.  Started instrument.  Azimuth seems to move freely but still getting motor errors.  RTS at 2035 GMT.

1/18/12

PAD

Last SDS collection was at 0700 GMT.  System argument out of range exception error is shown.  Az stall and Spec Error boxes are red.  Cleared error and exited program.  Rebooted computer.  At 1620 GMT, observed shadowband is not moving as program status indicates.  Motor Error:  MDrive send failed in Party Mode is shown.  This has been an ongoing error.  Notified instrument mentor.  RTS at 1337 GMT.

1/17/12

PAD

Shadowband is not in positions or moving as program status indicates.  Motor Error:  MDrive send failed in Party Mode is shown.  This has been an ongoing error.  Previous attempts have failed to find or correct problem.  Stopped restarted program but no change in status.  Notified instrument mentor. RTS at 1629 GMT.

1/10/12

PAD

Last SDS collection was at 0000 GMT.  Remotely accessed computer at 1306 GMT.  SASpect program is up but nothing is visible, black screen.  Stopped/restarted program but nothing started.  Only thing shown in status was a tilt angle warning.  Remotely rebooted computer and tried to start program again. Same results.  At 1533 GMT, shut down computer and instrument and powered up.  Zeroed azimuth and inner band and started program.  Program started correctly and shadowband is shading.  RTS at 1548 GMT.

1/9/12

PAD

Last SDS collection was at 0100 GMT.  At 1517 GMT, observed both motor error and Az stall boxes are sometimes red.  Shadowband is not doing what is indicated on Saspect program.  Stopped program and zeroed azimuth and inner band.  Started program with same results.  Stopped program and rebooted computer.  Started program after computer boot up with same results.  Troubleshooting at this time.  RTS at 1537 GMT.  At 1939 GMT, stopped instrument and removed cover.  Lubricated shadowband shaft at drive and driven ends with greaseless lubricant.  Reset mechanical sprocket alignment with shadowband horizontal positions.  Started program.  Shadowband and azimuth position are now in correct alignment.  RTS at 2039 GMT.

1/3/12

PAD

Last SDS collection was on 1/2/12 at 1900 GMT.  Saspect program is open but covered by screen saver.  Exited program and rebooted computer.  Started program but shadowband is not shading correctly.  Reset offset for Inner band and Azimuth.  Shadowband is not shading correctly.  RTS at 1606 GMT.

12/29/11

PAD

Last SDS collection was at 0200 GMT.  Was able to remotely access computer at 1320 GMT but computer was not responsive to commands.  At 1523 GMT computer was not responsive to commands or icons.  Shut down/restarted computer.  Set inner band offset and restarted program.  Shadowband is shading correctly.  RTS at 1538 GMT.

12/9/11

PAD

Last SDS collection was at 1800 GMT.  At 2310 GMT, remotely accessed computer and found SASpect program was shown but stopped.  Started program.  RTS at 2325 GMT.

12/5/11

PAD

Last SDS collection was on 12/4/11 at 0400 GMT.  Remotely accessed computer and found SASpect program was open but nothing shown except white box.  Stopped/restarted SASpect program.  Program started correctly.  At 1548 GMT, observed at computer SASpect program was not shown.  Computer is not responsive to ctl/alt/del to bring up applications, nor will reboot or shutdown option come up when start menu is selected.  Powercycled computer and started SASpect program after boot up.  Program started correctly but too cloudy to verify shading.  RTS at 1600 GMT.

11/22/11

PAD

SAS-He refrigerator is getting moisture inside.   Mentor has sent a replacement door that has been altered to allow dry air to flow into refrigerator.  R/R original door with replacement provided.  Connected air line from dry air system located in the Optical trailer.  Restarted program and had to adjust Azimuth offset so instrument would shade correctly.  RTS at 2156 GMT.

10/24/11

PAD

Shadowband has not been working correctly for some time.  Mentor sent a replacement Az drive motor and requests it be installed.  Removed Az drive motor and installed replacement.  Motor error was shown after instrument startup:  Motor axis Az iMdrive failed in Party Mode State=Alarm; Control variable=0.  Further troubleshooting found error is only shown when replacement motor is connected to comms cable.  Az and both shadowband motors share the same cable.  Mentor suspects a problem with replacement motor and requested motor be removed and sent back to him.  Removed motor and took to shipping dept.  Original Model MDI3CRD17A4-EQ, SN 127090035.  Replacement Model Model MDI3CRD17A4-EQ, SN 036100216.  RTS at 2038 GMT.

10/13/11

PAD

Shadowband is not doing what system status shows.  Mentor requests technician on site to assist in troubleshooting.  Assisted mentor remotely troubleshooting shadowband issue including swapping motors, moving shadowband to outer band motor, intermittent limit switch failure, etc.  Ended in inner and outer motors, motors A and C, back in original motor mounts.  Shadowband is being driven by motor A.  Still experiencing intermittent failure of one of the optical limit switches for motor A.  Mentor is trying to adjust software to only use one limit switch.  RTS at 1940 GMT.

10/12/11

PAD

Shadowband is not doing what system status shows.  Mentor requests technician on site to assist in troubleshooting.  Assisted mentor remotely troubleshooting shadowband issue including swapping motors, switching motor power wires between inner and outer motors, intermittent limit switch failure, etc.  Failed to isolate problem.  RTS at 2230 GMT.

10/11/11

PAD/KT

SAS-He shadowband has not been doing what system status indicates.  This issue has been intermittent but ongoing for last couple of weeks.  Previous troubleshooting has failed to isolate cause.  Mentor requests shadowband moved to outer drive motor and horizontal limits set on shadowband.  Moved shadowband as mentor requested according to instructions provided.  After changing shadowband, outer band motor fails to recognize one of the limit switches.  Troubleshooting found limit switch is lighting up when tripped indicating it is working correctly.  Mentor then requested inner and outer motors switched and shadowband moved back to inner band motor shaft, original location.  At 1900 GMT, disconnected, removed, swapped, and installed inner and outer motors.  Moved shadowband as requested to original location.  Contacted mentor after changes were made.  Mentor is currently working with software engineer trying to make necessary changes in order to make outer band motor work in inner band motor mounted position.  RTS at 2010 GMT.

10/5/11

PAD

Shadowband is not doing what is indicated on system status.  Shadowband is re-zeroing but is not shading diffuser.  Band 1 low and Motor error boxes are red.  Stopped program and rebooted computer.  Shadowband is now shading correctly.  RTS at 1516 GMT.

10/4/11

PAD

Shadowband is not doing what is indicated on system status.  Shadowband is re-zeroing but is not shading diffuser.  Stopped program and rebooted computer.  Did not put experiment into run continuous until 1455 GMT.  Shadowband is now shading correctly.  RTS at 1500 GMT.

10/3/11

PAD

Shadowband is not doing what software status is showing.  Shadowband is re-zeroing but is not shading diffuser.  Instrument is making a clicking noise when shadowband is horizontal.  Band 1 low, motor error, and Az stall boxes are red on monitor.  Stopped program and rebooted computer.  After restart, instrument is now operating correctly.  RTS at 1409 GMT.

9/29/11

PAD

Status shows instrument going through 19 measurement steps but instrument is not doing what status is showing.  Stopped SASpect program and rebooted computer.  After restart, adjusted Az offset to 16...   Shadowband is still not doing what instrument status shows.  Notified instrument mentor.  Troubleshooting at this time.  RTS at 1428 GMT.

9/28/11

PAD

Status shows instrument going through 19 measurement steps but instrument is not doing what status is showing.  Stopped SASpect program and rebooted computer.  After restart, adjusted Az offset to 16...  Instrument is now operating and shading correctly.  RTS at 1427 GMT.

9/26/11

PAD

Last SDS collection was on 9/25/11 at 1100 GMT.  Error -3 shown.  Cleared error but it kept reoccurring.  Used task manager to end SASpect program.  Rebooted computer and changed Az offset to 16...  Instrument is now shading correctly.  RTS at 1403 GMT.

9/19/11

PAD

Error shown:  ConnorSpec has encountered a problem and needs to close.  Cleared error message and SASpect program closed.  Rebooted computer and after startup, set azimuth offset to 16.####.  Instrument is now running and shadowband is correct.  RTS at 1538 GMT.

9/16/11

SPD

Error on screen, "ConnerSpec encountered a problem and needs to close".  GPS sync 71198.3 box was red.  Band 1 low 260.7 box was red.   Motor error box was red.   Clicked the don't send button on error page and SAS Spec closed. Restarted the SAS Spec program. RTS 1537 GMT.

9/15/11

PAD

Error shown:  ConnorSpec has encountered a problem and needs to close.  Printed screen and saved error message.  Rebooted computer and reset azimuth offset after restart.  Unable to tell if instrument is shading due to sky conditions.  RTS  at 1428 GMT.

9/14/11

PAD

Program shows instrument is taking measurements but shadowband is continually rezeroing.  Stopped/restarted program and entered correct azimuth offset.  Error was shown:  National Instruments.  DAQmx. DaqException.  Took screenshot of error and saved.  Program was not responsive to stop command.  Program shows instrument is on step 2 of 19, Acquiring Dark low limit.  Used task manager to stop program.  Rebooted computer.  After program restarted, stopped program and reset azimuth offset.  Instrument is now operating and shading is correct.  RTS at 1606 GMT.

9/8/11

PAD

Spectrometer was removed on 8/29/11 per mentor's request.  Mentor requests spectrometer placed back into service.   Mentor will attempt to run system with one spectrometer.  Installed spectrometer and connected all power and comms.  Powered up system and program started correctly with no errors shown as when spectrometer failed to power up on 8/29/11.  Notified mentor.  RTS at 2105 GMT.

8/29/11

PAD

Unit was shut down on 8/25/11 to allow freezer housing spectrometers to thaw.  Instrument would not restart correctly on 8/26/11.  Received T/S instructions from mentor via email.  At 1327 GMT, found spectrometer NIR256-1.7 was not powering up.  Troubleshooting found 5.2V DC input voltage at USB splitter that powers both spectrometers.  Removed spectrometer at Mentor's request and took to lab.  12V external power supplied to spectrometer will power it up but still unable to communicate via USB port.  Mentor suspects USB port failure on spectrometer.  Waiting on replacement at this time.  Model AvaSpec-NIR256-1.7, SN 1010110U1, WD66832.  RTS at 2000 GMT.

8/26/11

SPD

Cleaned freezer and replaced spectrometer restarted computer. SAS SPEC program will not run there are SEVERAL errors that pop up.  Trouble shooting still in progress. Notified Mentor via a voice mail, awaiting his response.

8/24/11

PAD

Shadowband is not shading correctly.  Azimuth position is off.  Initialized and reset Azimuth offset.  Shadowband is now shading correctly.  RTS at 1430 GMT.

8/22/11

PAD

Notified by mentor he suspects condensation in refrigerator housing SAS-He spectrometer.  Found significant ice buildup on heat sinks located in refrigerator.  Consulted with mentor and was asked to remove rack which spectrometer is attached from refrigerator and allow refrigerator to defrost overnight.  Shut down instrument and removed spectrometer as instructed.  Instrument will be restarted tomorrow.  RTS at 1952 GMT.

8/22/11

PAD

Notified by mentor he suspects condensation in refrigerator housing SAS-He spectrometer.  Found significant ice buildup on heat sinks located in refrigerator.  Consulted with mentor and was asked to remove rack which spectrometer is attached from refrigerator and allow refrigerator to defrost overnight.  Shut down instrument and removed spectrometer as instructed.  Instrument will be restarted tomorrow.  RTS at 1952 GMT.

8/12/11

PAD

Shadowband is not shading diffuser.  Azimuth position is not correct.  Used software to align instrument with sun and manually set offset.  Shadowband is now shading correctly.  RTS at 1347 GMT.

7/27/11

SD

Noticed  a RED Motor Error on display at 1600 GMT. Also noticed that the Temp Graph has unrealistic X and Y numbers.  Contacted mentor. He viewed the problems remotely and will work on them.

7/21/11

JEM

Shadowband is not shading correctly.  Adjusted Azimuth offset -4 deg.  Restarted program.  Shadowband is now shading correctly.  RTS at 1343 GMT.

7/5/11

PAD

Error shown:  ConnorSpec has encountered an error and needs to close.  Cleared error and SASpect program closed.  Restarted program.  RTS at 1226 GMT.

6/21/2011

JEM

The SAS-He has been experiencing issues with only operating properly during the morning hours.  The mentor indicates there may be a cable/connection issue inside the instrument housing.  Opened the instrument housing and was able to make the SAS-He fail at will by manipulating the cable that runs between the lower and upper units that controls the limit switches.  Using a meter found the troubled wire inside the cable.  Cut the blue wire at each end of the cable and spliced in a new wire to bypass the bad section.  After repairing the cable, re-zeroed the azimuth and the band.  Restarted the experiment @1545 GMT.  The SASA-He has returned to normal operation.  RTS @1600 GMT.

6/17/2011

JEM

The SAS is not operating.  Started the program but the band would not move.  The band offered no resistance when moved by hand.  Stopped the SAS-He program.  Removed the cover and inspected the band motor.  Found the different set screws that had fallen out and were lying in the bottom of the enclosure.  Put the set screws back where they belonged (flat side of shafts).  Restarted the program.  Instrument is working now, but alignment is not right.  Since the band does an auto zero and is still not aligning properly, I am going to notify the mentor and seek his advice on further action.  Did not RTS.

6/16/11

PAD

Shadowband has not been operating correctly for any 24 hour period for the last two weeks.  At 1504 GMT, observed shadowband past horizontal and not active.  Program shows step 19 of 19, inner band position shows 1,134,098 and counting.  Selected kill motion for inner band several times, band would try to move but very slowly.  Exited program, rebooted computer and selected run continuous experiment.  Shadowband is still trying to move but very slowly.  Very little resistance felt when manually moving shadowband.  Will remove instrument cover and TS when wind is lighter.  RTS at 1519 GMT.

6/13/11

PAD

Error shown:  ConnorSpec has encountered a problem and needs to close.  Printed error message and saved on desktop.  Started SASpect program.  Instrument did start up and zero shadowband.  Unable to verify correct shadowband alignment due to clouds.  RTS at 1427 GMT.

6/10/2011

JEM

The SAS-He is not operating correctly.  The shadowband is bouncing at it's limit and the band location reading is a very high negative number and continuing to drop.  Stopped the band motion.  Closed the SASPect program.  Rebooted the instrument PC.  After the PC and SASPect program came up.  Selected get position on the band.  The position provided was not accurate, as it was a large negative number.  Selected experiment from the menu then chose continuous.  The band position corrected itself and now the SAS-He is operating normally.  RTs @1509 GMT.

6/9/2011

JEM

The SAS-He is not operating correctly.  The shadowband is bouncing at it's limit and the band location reading is a very high negative number and continuing to drop.  Stopped the band motion.  Closed the SASPect program.  Rebooted the instrument PC.  After the PC and SASPect program came up.  Selected get position on the band.  The position provided was not accurate, as it was a large negative number.  Selected experiment from the menu then chose continuous.  The band position corrected itself and now the SAS-He is operating normally.  RTS @1428 GMT.

6/8/2011

JEM

The SAS-He is not operating correctly.  The shadowband is bouncing at its limit and the band location reading is a very high negative number and continuing to drop.  Stopped the band motion.  Closed the SASPect program.  Rebooted the instrument PC.  After the PC and SASPect program came up.  Selected get position on the band.  The position provided was not accurate, as it was a large negative number.  Selected experiment from the menu then chose continuous.  The band position corrected itself and now the SAS-He is operating normally.  The mentor knows about this problem but he is in India at the current time.  RTS @1545 GMT.

6/6/2011

JEM

The SAS-He is not operating correctly.  The shadowband is bouncing at its limit and the band location reading is a very high negative number and continuing to drop.  Rebooted the computer @1437 GMT.  After the computer booted up, ran the continuous experiment.  RTS @1445 GMT.

5/16/11

PAD

At 1436 GMT, observed shadowband is not shading diffuser.  Notified mentor via voicemail.  Mentor called at 1620 GMT.  Worked with mentor who was remotely accessing computer to align instrument shadowband and Azimuth.  Instrument is now shading correctly.  RTS at 1730 GMT.

5/5/11

PAD

Saspect program is up but screen saver is covering it.  Instrument is not active.  Printed screen and saved.  Tried to close Saspect program but it will not respond.  National Instrument error message was then shown.  Printed screen of error message and saved.  Used task manager to end program.  Rebooted computer and set inner band zero.  Tried to set azimuth zero but it would not respond.  Left voicemail with mentor.  RTS at 1436 GMT.

4/21/11

PAD

Unable to RA computer.  At 1327 error message shown:  Microsoft. NET Framework.  Index out of range.  Printed screen and saved error message on desktop.  Started SASpect program and zeroed inner band.  Started Continuous Experiment.  Appears shadowband is working correctly but Azimuth is not in correct position.  Left voicemail with mentor.  RTS at 1342 GMT.

3/29/11

CAM

SAS-He has been writing SAS-Ze data files since being restarted after encountering 2 software errors on 3/28.  Mentor suspects changes he made last week perhaps were not retained after the software crashed.  Stopped SASpect program at 1747 GMT.  Changed the instrument type from Ze to He in the configuration file as per mentor’s instructions.  Rebooted computer and started program in continuous mode.  Confirmed configuration and that proper data files were being written.  Moved all SAS-Ze data files out of the data directory and saved in SasData-20110329.  RTS at 1756 GMT.

3/28/11

PAD

Last SDS collection was on 3/26/11 at 0208 GMT.  Error shown:  ConnorSpec has encountered a problem and needs to close.  There is a National Instrument error behind Connor Spec error.  Printed screen of error message and saved on desktop.  Cleared error message and both errors and SAS program closed out.  Started SASHE program and notified mentor.  RTS at 1356 GMT.

3/17/11

PAD

At 1336 GMT, observed shadowband is not shading diffuser.  Left voicemail with mentor.  At 1710 GMT, worked with mentor via phone to troubleshoot.  Shadowband would move but very slowly.  Shadowband is not slipping on shaft and limit stops can be detected.  Found limit for shadowband was set at 3…. And -3…  Changed settings to 90 and -90 as instructed by mentor.   Shadowband will now move at correct speed.  Aligned shadowband and instrument axis as per mentor’s instruction and restarted program.  National Instrument errors were shown after starting program.  Stopped program, rebooted computer and started program.  Error was shown again.  Cleared error and program started.  Shadowband is now shading correctly.  RTS at 1830 GMT.

3/11/11

Mentor

Instrument installed by Connor Flynn and Dan Nelson, 3/9-3/11.